OpenStudy (jdubh1411):

What is 6! over 3! ?

OpenStudy (jtvatsim):

Remember that 6! = 6 x 5 x 4 x 3 x 2 x 1 and 3! = 3 x 2 x 1.

OpenStudy (jtvatsim):

If we divide these two numbers (or as you say, 6! over 3!) we have \[\frac{6!}{3!} = \frac{6 \times 5 \times 4 \times 3 \times 2 \times 1}{3\times 2 \times 1}\] Simplify the fraction and you will be done!